OpenStudy (natasha18):

Can anyone teach me some basic tips?? I'm just curious lol

whitemonsterbunny17 (whitemonsterbunny17):

Yes, I can teach you.

whitemonsterbunny17 (whitemonsterbunny17):

The very first thing that you must know is to always start each command with `\(` and then end with `\)`

whitemonsterbunny17 (whitemonsterbunny17):

Here is an example code: `\(\Large\cal\color{blueviolet}{Hello!}\)`

OpenStudy (natasha18):

Okay. I've been looking around. I know a basic way to do one, but every time I post it, it just stays in code form..???

OpenStudy (natasha18):

@Whitemonsterbunny17

whitemonsterbunny17 (whitemonsterbunny17):

Can you post what you've got so far so that I can look at it and identify the error?

OpenStudy (natasha18):

@Whitemonsterbunny17 \(\color{red}{Natasha}\)

OpenStudy (natasha18):

OOOooooh ok. I think I know what I messed up on.

OpenStudy (natasha18):

I was using the wrong slash (/) XD

whitemonsterbunny17 (whitemonsterbunny17):

Ah, yeah, that'll do it! Sometimes I make that mistake too, don't worry. :p

whitemonsterbunny17 (whitemonsterbunny17):

So, now we can experiment with some different sizes and fonts. \(\Large\rm\color{lime}{~~~~~~Sizes:}\) ` \tiny ` = \(\tiny\rm\color{red}{tiny}\) ` \small ` = \(\small\rm\color{orange}{small}\) ` \normalsize ` = \(\normalsize\rm\color{gold}{normalsize}\) ` \large ` = \(\large\rm\color{lime}{large}\) ` \Large ` = \(\Large\rm\color{cyan}{Large}\) ` \LARGE ` = \(\LARGE\rm\color{cornflowerblue}{LARGE}\) ` \huge ` = \(\huge\rm\color{blueviolet}{huge}\)

OpenStudy (natasha18):

@Whitemonsterbunny17 Woooow that's awesome! :P

whitemonsterbunny17 (whitemonsterbunny17):

Thanks! :) The sizes are the second command to be typed (after your font name, which we will cover next). For example: `\(\rm\Large\color{blueviolet}{Whitemonsterbunny17}\)` ^

OpenStudy (natasha18):

@Whitemonsterbunny17 Okay. And thanks a lot for the help! :)

whitemonsterbunny17 (whitemonsterbunny17):

You're welcome! Here are the fonts:

whitemonsterbunny17 (whitemonsterbunny17):

\(\Large\bf\color{blueviolet}{~~~~~~~~Fonts:}\) ` \bf ` = \(\Large\bf\color{red}{Example}\) ` \rm ` = \(\Large\rm\color{orange}{Example}\) ` \tt ` = \(\Large\tt\color{gold}{Example}\) ` \it ` = \(\Large\it\color{lime}{Example}\) ` \cal ` = \(\Large\cal\color{cyan}{Example}\) ` \frak ` = \(\Large\frak\color{blue}{Example}\) ` \mathbb ` = \(\Large\mathbb{\color{blueviolet}{EXAMPLE}}\) \(\normalsize\rm\color{red}{Note:~Mathbb~is~a~special~case,~instructions~are~below:}\) \(\normalsize\rm\color{magenta}{A~bracket~must~be~placed~after~the~fontname~and~before~the~end~command.}\) \(\normalsize\rm\color{magenta}{Example:}\) `\(\Large\mathbb{\color{blueviolet}{EXAMPLE}}\)` \(\normalsize\rm\color{magenta}{Also~note~that~mathbb~only~changes~capitalized~letters~and~numbers.}\)

OpenStudy (natasha18):

It gets even \(\large\mathbb{\color{purple}{COOLER}}\) !! XD

whitemonsterbunny17 (whitemonsterbunny17):

Wow! That looks awesome! :) You're doing great! Now, if you want to write a sentence with multiple words, just use ~ to insert a normal space.

whitemonsterbunny17 (whitemonsterbunny17):

Example: `\(\large\rm\color{blueviolet}{Hello,~I~am~bunny!}\)` Which makes: \(\large\rm\color{blueviolet}{Hello,~I~am~bunny!}\)

OpenStudy (natasha18):

Sweet!! Does this work if you're typing in a word processor like Word and stuff?

whitemonsterbunny17 (whitemonsterbunny17):

No, it only works here and on a few other sources that have a LaTeX render.

whitemonsterbunny17 (whitemonsterbunny17):

For here, it's really helpful to use when teaching, as it makes it easier to highlight certain words or give a more visual teaching experience.

OpenStudy (natasha18):

Okay, gotcha :)

whitemonsterbunny17 (whitemonsterbunny17):

You can go ahead and try experimenting and practicing if you'd like. Practice makes perfect! :)

OpenStudy (natasha18):

@Whitemonsterbunny17 Okay I will. Again, thanks for all the help ^-^